What volume of 5.0 M HNO3 is required to neutralize 2500 cm3 of a 2.0 M NaOH solution
1
Expert's answer
2015-02-05T09:33:34-0500
The chemical equation for the reaction is:
NaOH+ HNO3 = NaNO3 + H2O
According to the equation the number of moles of the acid is equal to the number of moles of the base.
n(NaOH)= n(HNO3)
n(mol)= C(M)∙V(L)
C(NaOH)∙V(NaOH)= C(HNO3)∙V(HNO3)
2.0M∙ 2.5L = 5.0M∙V(HNO3)
V(HNO3)=(2.0∙2.5)/5.0 = 5.0/5.0 = 1 L
Answer: V(HNO3)=1 L